# A Dimca-Greuel type inequality for foliations

### v1

@@ 3. Proof of Theorem A
Let $C$ be an irreducible plane curve of multiplicity $\nu_{p}(C)=n$. Denote by $\tilde{C}$ the strict transform of $C$ by a blow-up and $\bar{C}$ the normalization of $C$. We have natural morphisms
(8) $\pi:\bar{C}\buildrel\rho\over{\longrightarrow}\tilde{C}\buildrel\sigma\over{\longrightarrow}C.$
Consider $(C,p)\buildrel i\over{\longrightarrow}(\mathcal{C},x_{0})\buildrel\phi\over{\longrightarrow}(T,t_{0})$ the miniversal deformation of $C$. We extend to $\mathcal{C}$ the morphisms in (8):
(9) $\pi:\bar{\mathcal{C}}\buildrel\rho\over{\longrightarrow}\tilde{\mathcal{C}}\buildrel\sigma\over{\longrightarrow}\mathcal{C}.$
We have two other natural morphisms $\tilde{C}\buildrel\tilde{\phi}\over{\longrightarrow}T$ and $\bar{C}\buildrel\bar{\phi}\over{\longrightarrow}T$ and its respectively extensions to the deformation, that is $\tilde{\mathcal{C}}\buildrel\tilde{\phi}\over{\longrightarrow}T$ and $\bar{\mathcal{C}}\buildrel\bar{\phi}\over{\longrightarrow}T$.
Let $\Omega_{\mathcal{C}/T}$ (respectively, $\Omega_{\tilde{\mathcal{C}}/T}$, respectively, $\Omega_{\bar{\mathcal{C}}/T}$ ) denote the $\mathcal{O}_{\mathcal{C}}$-module of relative Kähler differentials of $\mathcal{C}$ over $T$ (respectively, the $\mathcal{O}_{\tilde{\mathcal{C}}}$-module of relative Kähler differentials of $\tilde{\mathcal{C}}$ over $T$, respectively the $\mathcal{O}_{\bar{\mathcal{C}}}$-module of relative Kähler differentials of $\bar{\mathcal{C}}$ over $T$). Denote by $G:=\rho^{*}\Omega_{\tilde{\mathcal{C}}/T}/\pi^{*}\Omega_{{\mathcal{C}}/T}$. If $C$ is singular then $G$ is not zero and $\bar{\phi}_{*}G$ is a finite $\mathcal{O}_{T}$-module and after Wang ([15]) we put $\mathcal{D}(t)=\hbox{\rm length}_{\mathcal{O}_{T,t}}((\bar{\phi}_{*}G)_{t})$ and $\alpha:=\min_{t\in T}\mathcal{D}(t)$. By [15, Claim 4.4] we have
(10) $\alpha>\dfrac{\nu_{p}(C)(\nu_{p}(C)-1)}{4},$
for $\nu_{p}(C)\geq 2$.
We will need the following lemma.
@@
Let $\mathcal{F}$ be a germ of a singular foliation on $(\mathbb{C}^{2},p)$, and $B$ be a separatrix of $\mathcal{F}$ at $p$. If $\tilde{\mathcal{F}}$ (respectively $\tilde{B}$) is the strict transform of $\mathcal{F}$ (respectively of $B$) at the point $q$, then
(11) $3\mu_{q}(\tilde{\mathcal{F}},\tilde{B})-4\tau_{q}(\tilde{\mathcal{F}},\tilde{B})+GSV_{q}(\tilde{\mathcal{F}},\tilde{B})\geq 3\mu_{p}(\mathcal{F},B)-4\tau_{p}(\mathcal{F},B)+GSV_{p}(\mathcal{F},B).$
Moreover, the equality holds if and only if $B$ is smooth.
@@
where
(13) $m_{p}(\mathcal{F})=\left\{\begin{array}[]{ll}\nu_{p}(\mathcal{F})+1&\hbox{\rm if$\sigma$ is dicritical}\\ \nu_{p}(\mathcal{F})&\hbox{\rm if $\sigma$ is nondicritical}.\\ \end{array}\right.$
On the other hand by the behaviour under blow-up of the GSV index [3, p.30] we get
(14) $GSV_{p}(\tilde{\mathcal{F}},\tilde{B})=GSV_{p}(\mathcal{F},B)+\nu_{p}(B)(\nu_{p}(B)-m_{p}(\mathcal{F})),$
and by [7, Proposition 6.2] we have
@@
After [15, Equation (4)] we get:
(16) $\tau_{p}(B)-\tau_{q}(\tilde{B})=\frac{\nu_{p}(B)(\nu_{p}(B)-1)}{2}+\alpha,$
where $\alpha$ was defined in (10). From (15), (14), and (16) we deduce
(17) $\tau_{q}(\tilde{\mathcal{F}},\tilde{B})=\tau_{p}(\mathcal{F},B)+\dfrac{\nu_{p}(B)[\nu_{p}(B)+1-2m_{p}(\mathcal{F})]}{2}+\alpha.$
@@
(18) $3\mu_{q}(\tilde{\mathcal{F}},\tilde{B})-4\tau_{q}(\tilde{\mathcal{F}},\tilde{B})+GSV_{q}(\tilde{\mathcal{F}},\tilde{B})$ $=$ $3\mu_{p}(\mathcal{F},B)-4\tau_{p}(\mathcal{F},B)+GSV_{p}(\mathcal{F},B)$
-- $\nu_{p}(B)(\nu_{p}(B)-1)+4\alpha.$
Hence, if $B$ is a singular curve then the lemma follows by (10).
Suppose now that $B$ is a non-singular curve. By a change of coordinates, if necessary, we can suppose that $B$ is given by $x=0$. Since $B$ is a separatrix of $\mathcal{F}:\omega=Pdx+Qdy$ then $\omega\wedge dx=Qdx\wedge dy$. Hence $Q=xh$ for some formal power series $h(x,y)\in\mathbb{C}[[x,y]]$ such that $h(0,y)\not=0$ and $\mathcal{F}:\omega=P(x,y)dx+xh(x,y)dy$. Since $\gamma(t)=(0,t)$ is a parametrization of $B$, then $\mu_{p}(\mathcal{F},B)=\text{ord}_{t}P(0,t)$ and
(19) $\tau_{p}(\mathcal{F},B)=\dim_{\mathbb{C}}\mathbb{C}[[x,y]]/(P,xh,x)=\dim_{\mathbb{C}}\mathbb{C}[[x,y]]/(P,x)=\mu_{p}(\mathcal{F},B).$
Moreover by (15) we have the equality $GSV_{p}(\mathcal{F},B)=\tau_{p}(\mathcal{F},B)$. Thus if $\nu_{p}(B)=1$ then $\nu_{p}(\tilde{B})=1$, $\mu_{p}(\mathcal{F},B)=\tau_{p}(\mathcal{F},B)=GSV_{p}(\mathcal{F},B)$ and $\mu_{p}(\tilde{\mathcal{F}},\tilde{B})=\tau_{p}(\tilde{\mathcal{F}},\tilde{B})=GSV_{p}(\tilde{\mathcal{F}},\tilde{B})$, so (11) becomes an equality.
Moreover, if (11) is an equality, then $B$ is necessarily smooth. Indeed if $B$ were singular, by (10) the number $-\nu_{p}(B)(\nu_{p}(B)-1)+4\alpha$ is positive and from the equality (18) we have that (11) is a strict inequality that is a contradiction. Hence the lemma follows. ∎
@@
Let $\mathcal{F}$ be a germ of a singular foliation on $(\mathbb{C}^{2},p)$, and $B$ be a separatrix of $\mathcal{F}$ at $p$. If $\tilde{\mathcal{F}}$ (respectively $\tilde{B}$) is the strict transform of $\mathcal{F}$ (respectively of $B$) at the point $q$ after a composition of $\ell$ blows-up then
(20) $3\mu_{q}(\tilde{\mathcal{F}},\tilde{B})-4\tau_{q}(\tilde{\mathcal{F}},\tilde{B})+GSV_{q}(\tilde{\mathcal{F}},\tilde{B})\geq 3\mu_{p}(\mathcal{F},B)-4\tau_{p}(\mathcal{F},B)+GSV_{p}(\mathcal{F},B).$
Proof.
It follows applying Lemma 3.1 $\ell$ times. ∎
Remark 3.3.
By definition, for any branch $B$ we have $\tau_{p}(B)\leq\mu_{p}(B)$, so after [7, Proposition 6.2] and [7, Propositon 5.7] we get $\tau_{p}(\mathcal{F},B)-GSV_{p}(\mathcal{F},B)\leq\mu_{p}(\mathcal{F},B)-GSV_{p}(\mathcal{F},B)$ and
(21) $\tau_{p}(\mathcal{F},B)\leq\mu(\mathcal{F},B).$
In addition, by (19) if $B$ is smooth then $\tau_{p}(\mathcal{F},B)=\mu(\mathcal{F},B)$.
Remark 3.4.
If $\mathcal{F}$ is a non-dicritical second type foliation at $p\in\mathbb{C}^{2}$, and $B$ is the only separatrix of $\mathcal{F}$, then $\mathcal{F}$ is a generalized curve foliation. In fact, assume by contradiction that $\mathcal{F}$ admits a saddle-node singularity in its reduction process with strong separatrix on the exceptional divisor. Without lost of generality, we can suppose that $\mathcal{F}$ is a reduced foliation after a blow-up $\sigma$ at $p$. Since $\mathcal{F}$ is of second type, after a blow-up, the strong separatrix of the strict transform $\tilde{\mathcal{F}}$ of $\mathcal{F}$ through $q\in\sigma^{-1}(p)$ is contained in $E=\sigma^{-1}(p)$, thus the Camacho-Sad index is $CS_{q}(\tilde{\mathcal{F}},E)=0$, but it is a contradiction with Camacho-Sad’s formula
$CS_{q}(\tilde{\mathcal{F}},E)=E\cdot E=-1.$
Proof of Theorem A: If $B$ is a non-singular curve then, by Remark 3.3, $\mu_{p}(\mathcal{F},B)=\tau_{p}(\mathcal{F},B)$ and the theorem follows.
Let us suppose now that $B$ is a singular separatrix, that is $\nu_{p}(B)\geq 2$. After Corollary 3.2 we can suppose, without lost of generality that the number of blows-up $n$ needed in the process of the reduction of the singularity of $B$ is one. In particular, after a blow-up $\sigma$ the strict transform $\tilde{B}$ of $B$ is smooth, that is the algebraic multiplicity of $\tilde{B}$ is one.
Applying Lemma 3.1 and after (14) we get
(22) $[3\mu_{q}(\tilde{\mathcal{F}},\tilde{B})-4\tau_{q}(\tilde{\mathcal{F}},\tilde{B})]-[3\mu_{p}(\mathcal{F},B)-4\tau_{p}(\mathcal{F},B)]>\nu_{p}(B)(m_{p}(\mathcal{F})-\nu_{p}(B)).$
Since $\tilde{B}$ is smooth, from (22) and the equality $\tau_{q}(\tilde{\mathcal{F}},\tilde{B})=\mu_{q}(\tilde{\mathcal{F}},\tilde{B})$ (see Remark 3.3) we obtain
(23) $-\tau_{q}(\tilde{\mathcal{F}},\tilde{B})-[3\mu_{p}(\mathcal{F},B)-4\tau_{p}(\mathcal{F},B)]>\nu_{p}(B)(m_{p}(\mathcal{F})-\nu_{p}(B)).$
Moreover, by the definition of $m_{p}(\mathcal{F})$ and after [11, Proposition 3.3], if $\mathcal{B}=B+\mathcal{B}^{\prime}$ is a balanced divisor of separatrices of $\mathcal{F}$ (for some divisor $\mathcal{B}^{\prime}$, if $B$ is the only separatrix of $\mathcal{F}$ we consider for $\mathcal{B}^{\prime}$ the empty divisor) then we have
$m_{p}(\mathcal{F})-\nu_{p}(B)=\left\{\begin{array}[]{ll}\nu_{p}(\mathcal{B}^{\prime})-1+\xi_{p}(\mathcal{F})&\hbox{\rm if $\sigma$ is non-dicritical}\\ \nu_{p}(\mathcal{B}^{\prime})+\xi_{p}(\mathcal{F})&\hbox{\rm if $\sigma$ is dicritical},\end{array}\right.$
where $\xi_{p}(\mathcal{F})$ is the tangency excess of $\mathcal{F}$ (see for example [7, Definition 2.3]). So $m_{p}(\mathcal{F})-\nu_{p}(B)\geq 0$ except in the case that $\mathcal{F}$ is a foliation of the second type ($\xi_{p}(\mathcal{F})=0$) and $B$ is its only separatrix ($\nu_{p}(\mathcal{B}^{\prime})=0$) in which case we get $m_{p}(\mathcal{F})-\nu_{p}(B)=-1$. In the case $m_{p}(\mathcal{F})-\nu_{p}(B)\geq 0$, from (23) we obtain $03\mu_{p}(\mathcal{F},B)-4\tau_{p}(\mathcal{F},B)<$ 0 and the theorem follows. Otherwise, assume that $\mathcal{F}$ is of the second type and that $B$ is its only separatrix. By Remark 3.4, $\mathcal{F}$ is a generalized curve foliation, so $GSV_{p}(\mathcal{F},B)=0$. Moreover, since $\tilde{B}$ is smooth then the left-term of (11) is zero, so the theorem follows by Lemma (3.1) (since $B$ is singular the inequality (11) is strict) and Remark 3.3 ($\tilde{B}$ is smooth). $\Box$
From Theorem A we deduce the Dimca-Greuel bound for the case of irreducible plane curve germs, proved simultaneously in [1], [10] and [15]:
Corollary 3.5.
Let $f(x,y)\in\mathbb{C}[[x,y]]$ an irreducible power series. Put $C:f(x,y)=0$ the branch defined by $f$. Then
@@
It is a consequence of Remark 2.1 and of applying Theorem A to foliation $\mathcal{F}:df=0$. ∎
Remark 3.6.
In [15] Wang proved that the number $3\mu_{p}(B)-4\tau_{p}(B)$ is a non-negative number, non-decreasing under blowups and strictly increasing unless the curve $B$ is non-singular. The result of Wang is an special case of Lemma 3.1, taking $\mathcal{F}=df$ where $B$ is given by $f(x,y)=0$. Lemma 3.1 and Theorem A generalizes this for the number $3\mu_{p}(\mathcal{F},B)-4\tau_{p}(\mathcal{F},B)$.
Now, if in (1), we change $\mu_{p}(\mathcal{F})$ to the multiplicity of the singular foliation $\mathcal{F}$ along a reduced curve $C$, $\mu_{p}(\mathcal{F},C)$ (see Section 2), and we put in the denominator the sum $T_{p}(\mathcal{F},C):=\sum_{B}\tau_{p}(\mathcal{F},B)$ of Tjurina numbers of $\mathcal{F}$ along the irreducible components of $C$, then we get the following inequality of the Dimca-Greuel type for foliations:
@@
Let $\mathcal{F}$ be a germ of a singular foliation at $(\mathbb{C}^{2},p)$. Let $\mathcal{B}=\mathcal{B}_{0}-\mathcal{B}_{\infty}$ be a balanced divisor of separatrices for $\mathcal{F}$ at $p$. Then
(24) $\frac{\mu_{p}(\mathcal{F},\mathcal{B})}{T_{p}(\mathcal{F},\mathcal{B}_{0})}<\frac{4}{3}.$
@@
By (6) we have $\frac{\mu_{p}(\mathcal{F},\mathcal{B})}{T_{p}(\mathcal{F},\mathcal{B}_{0})}\leq\frac{\mu_{p}(\mathcal{F},\mathcal{B}_{0})}{T_{p}(\mathcal{F},\mathcal{B}_{0})}$. Hence we will prove that
(25) $\frac{\mu_{p}(\mathcal{F},\mathcal{B}_{0})}{T_{p}(\mathcal{F},\mathcal{B}_{0})}<\frac{4}{3}.$
Suppose that $\mathcal{B}_{0}=\sum_{j=1}^{\ell}B_{j}$, where each $B_{j}$ is a separatrix for $\mathcal{F}$. It follows from Theorem A that
@@
$\frac{\mu_{p}(\mathcal{F},\mathcal{B}_{0})}{T_{p}(\mathcal{F},\mathcal{B}_{0})-\ell+1}<\frac{4}{3}.$
Since $\ell\geq 1$, we obtain $\frac{\mu_{p}(\mathcal{F},\mathcal{B}_{0})}{T_{p}(\mathcal{F},\mathcal{B}_{0})}\leq\frac{\mu_{p}(\mathcal{F},\mathcal{B}_{0})}{T_{p}(\mathcal{F},\mathcal{B}_{0})-\ell+1}<\frac{4}{3}.$

### v2

@@ 4. Proof of Theorem A
First of all, we will study the behaviour of $3\mu_{p}(\mathcal{F},B)-4\tau_{p}(\mathcal{F},B)$ under a blow-up $\sigma$ at $p\in\mathbb{C}^{2}$. For this, we consider $\alpha_{0}$ as in (14) for $j=0$.
We have the following lemma.
@@
Let $\mathcal{F}$ be a germ of a singular holomorphic foliation on $(\mathbb{C}^{2},p)$, and $B\in\text{Sep}_{p}(\mathcal{F})$. If $\tilde{\mathcal{F}}$ (respectively $\tilde{B}$) is the strict transform of $\mathcal{F}$ (respectively of $B$) by a blow-up $\sigma$ at $p$ then
$3\mu_{q}(\tilde{\mathcal{F}},\tilde{B})-4\tau_{q}(\tilde{\mathcal{F}},\tilde{B})=3\mu_{p}(\mathcal{F},B)-4\tau_{p}(\mathcal{F},B)-\nu_{p}(B)(2\nu_{p}(B)-m_{p}(\mathcal{F})-1)+4\alpha_{0},$
where $q\in\tilde{B}\cap\sigma^{-1}(p)$, and $\alpha_{0}$ is as in (14).
@@
where $m_{p}(\mathcal{F})$ was introduced in (17).
On the other hand by the behaviour under blow-up of the GSV index (see [4, p.30]) we get
(23) $GSV_{q}(\tilde{\mathcal{F}},\tilde{B})=GSV_{p}(\mathcal{F},B)+\nu_{p}(B)(\nu_{p}(B)-m_{p}(\mathcal{F})),$
and after [8, Proposition 6.2] we have
@@
From [19, Equation (4)] we obtain
(25) $\tau_{p}(B)-\tau_{q}(\tilde{B})=\frac{\nu_{p}(B)(\nu_{p}(B)-1)}{2}+\alpha_{0},$
where $\alpha_{0}$ was defined in (14). From (24), (23), and (25) we deduce
(26) $\tau_{q}(\tilde{\mathcal{F}},\tilde{B})=\tau_{p}(\mathcal{F},B)+\dfrac{\nu_{p}(B)[\nu_{p}(B)+1-2m_{p}(\mathcal{F})]}{2}-\alpha_{0}.$
@@
$3\mu_{q}(\tilde{\mathcal{F}},\tilde{B})-4\tau_{q}(\tilde{\mathcal{F}},\tilde{B})=3\mu_{p}(\mathcal{F},B)-4\tau_{p}(\mathcal{F},B)-\nu_{p}(B)(2\nu_{p}(B)-m_{p}(\mathcal{F})-1)+4\alpha_{0}.$
Remark 4.2.
If $\mathcal{F}$ is a reduced foliation at $p\in\mathbb{C}^{2}$, and $B$ is a smooth separatrix through $p$, then, by Remark 2.1, $GSV_{p}(\mathcal{F},B)=\mu_{p}(\mathcal{F}.B)=\tau_{p}(\mathcal{F},B)$. Therefore, according to [3, p. 538], we have $\tau_{p}(\mathcal{F},B)=1$, if $p$ is a non-degenerate singularity, or if $p$ is a saddle-node singularity and $B$ is the strong separatrix; and $\tau_{p}(\mathcal{F},B)=k+1$, if $p$ is a saddle-node singularity and $B$ is the weak separatrix. Here $k\in\mathbb{Z}^{+}$ is a formal invariant in the case that $p$ is a saddle-node singularity of $\mathcal{F}$. In any cases, $\tau_{p}(\mathcal{F},B)>0$.
We will now prove the main theorem of this paper.
4.1. Proof of Theorem A
Let $\mathcal{F}$ be a holomorphic foliation on $(\mathbb{C}^{2},p)$, and let $B$ be a separatrix of $\mathcal{F}$ at $p$. Then, on a minimal reduction of singularities $\pi:(\tilde{X},\mathcal{D})\to(\mathbb{C}^{2},p)$ of $\mathcal{F}$ we marked the reduction process of the pair $(\mathcal{F},B)$:
(27) $(\mathcal{F}^{(N)},B^{(N)})\overset{\sigma_{N}}{\longrightarrow}\cdots\overset{\sigma_{3}}{\longrightarrow}(\mathcal{F}^{(2)},B^{(2)})\overset{\sigma_{2}}{\longrightarrow}(\mathcal{F}^{(1)},B^{(1)})\overset{\sigma_{1}}{\longrightarrow}(\mathcal{F}^{(0)},B^{(0)})=(\mathcal{F},B),$
where $\pi=\sigma_{1}\circ\ldots\circ\sigma_{N}$, $B^{(j+1)}$ denotes the strict transform of $B^{(j)}$ thought $q^{(j)}$ for each $j=0,\ldots,N-1$, $B^{(N)}$ coincide with the normalization of $B$ (or a strict transform of it), $\mathcal{F}^{(i+1)}$ denotes the strict transform of $\mathcal{F}^{(j)}$, and $\mathcal{F}^{(N)}$ is a foliation with a reduced singularity at $q^{(N)}\in B^{(N)}\cap\mathcal{D}$. Put $\mu^{(i)}=\mu_{q^{(i)}}(\mathcal{F}^{(i)},B^{(i)})$ and $\tau^{(i)}:=\tau_{q^{(i)}}(\mathcal{F}^{(i)},B^{(i)})$, for any $i\in\{0,\ldots,N\}$.
Applying Lemma 4.1 to each blow-up $(\mathcal{F}^{(i+1)},B^{(i+1)})\overset{\sigma_{i+1}}{\longrightarrow}(\mathcal{F}^{(i)},B^{(i)})$, we get
$3\mu^{(i)}-4\tau^{(i)}=3\mu^{(i+1)}-4\tau^{(i+1)}+\left[\nu_{q^{(i)}}(B^{(i)})(2\nu_{q^{(i)}}(B^{(i)})-m_{q^{(i)}}(\mathcal{F}^{(i)})-1)-4\alpha_{i}\right]$
In particular
$3\mu_{p}(\mathcal{F},B)-4\tau_{p}(\mathcal{F},B)$ $=$ $3\mu^{(N)}-4\tau^{(N)}$
$+$ $\sum_{j=0}^{N-1}\left[\nu_{q^{(j)}}(B^{(j)})(2\nu_{q^{(j)}}(B^{(j)})-m_{q^{(j)}}(\mathcal{F}^{(j)})-1)-4\alpha_{j}\right],$
$=$ $3\mu^{(N)}-4\tau^{(N)}+R_{p}(\mathcal{F},B),$
where the last equality follows from (15).
We have two possibilities: either $B\in\text{ISO}_{p}(\mathcal{F},B)$ or $B\in\text{Dic}_{p}(\mathcal{F})$. If $B\in\text{ISO}_{p}(\mathcal{F},B)$, there exists a component $D\subset\mathcal{D}$ with $D$ non-dicritical such that $\mathcal{F}^{(N)}$ has a reduced singularity at $q^{(N)}\in\mathcal{D}\cap B^{(N)}$. In this case, we have $\mu^{(N)}=\tau^{(N)}$ by Remark 4.2, and thus
$3\mu_{p}(\mathcal{F},B)-4\tau_{p}(\mathcal{F},B)=-\tau^{(N)}+R_{p}(\mathcal{F},B)=\text{DG}_{p}(\mathcal{F},B).$
If $B\in\text{Dic}_{p}(\mathcal{F})$, then there exists a component $D\subset\mathcal{D}$ with $D$ dicritical such that $B^{(N)}$ is transversal to $D$ at $q^{(N)}$, and $\mathcal{F}^{(N)}$ is a regular foliation at $q^{(N)}$. In this case, we get
$3\mu_{p}(\mathcal{F},B)-4\tau_{p}(\mathcal{F},B)=R_{p}(\mathcal{F},B)=\text{DG}_{p}(\mathcal{F},B).$
In both cases, the hypothesis $\text{DG}_{p}(\mathcal{F},B)<0$ implies $03\mu_{p}(\mathcal{F},B)-4\tau_{p}(\mathcal{F},B)<$ 0 and the theorem follows. $\Box$
The following example illustrates that the hypothesis $\text{DG}_{p}(\mathcal{F},B)<0$ in Theorem A cannot be removed.
Example 4.3.
Let $\mathcal{F}_{m}$ be the germ of holomorphic foliation at $(\mathbb{C}^{2},0)$ defined by
$\omega_{m}=\left((2m+1)yx^{m+1}+my^{m+2}\right)dx+\left((1-m)xy^{m+1}-2mx^{m+2}\right)dy,$
where $m\in\mathbb{Z}_{>0}$. Let $B:x^{2m+1}+x^{m}y^{m+1}+y^{2m}=0$. We have that $B$ is a separatrix of $\mathcal{F}$ at $0\in\mathbb{C}^{2}$ such that
$GSV_{0}(\mathcal{F}_{m},B)=\left\{\begin{array}[]{cl}3&\text{if $m=1$}\\ -2m^{2}+4m+1&\text{ if $m>1$},\end{array}\right.$
$\mu_{0}(B)=2m(2m-1)\,\,\,\,\,\,\,\,\,\,m\geq 1,$
$\tau_{0}(B)=\left\{\begin{array}[]{cl}2&\text{if $m=1$}\\ 3m^{2}&\text{ if $m>1$},\end{array}\right.$
$\mu_{0}(\mathcal{F}_{m},B)=\left\{\begin{array}[]{cl}5&\text{ if $m=1$}\\ 2m^{2}+2m+1&\text{ if $m>1$}\end{array}\right.$
and
$\tau_{0}(\mathcal{F}_{m},B)=\left\{\begin{array}[]{cl}5&\text{ if $m=1$}\\ m^{2}+4m+1&\text{if $m>1$}.\end{array}\right.$
A straightforward calculation reveals that $\frac{\mu_{0}(\mathcal{F}_{m},B)}{\tau_{0}(\mathcal{F}_{m},B)}<\frac{4}{3}$ for $60<m<$ 6, and
$\lim_{m\to+\infty}\frac{\mu_{0}(\mathcal{F}_{m},B)}{\tau_{0}(\mathcal{F}_{m},B)}=2.$
On the other hand, we compute $\text{DG}_{0}(\mathcal{F}_{m},B)$ for $m\geq 1$. Note that $\nu_{0}(\mathcal{F}_{m})=m+2$, and $\nu_{0}(B)=2m$, and using [19, equation 4] we get
$\alpha_{0}=\left\{\begin{array}[]{cl}1&\text{ if $m=1$}\\ m^{2}+m&\text{ if $m>1$}\end{array}\right.$
After a blow-up at $0\in\mathbb{C}^{2}$, the strict transform of $B$ is a smooth curve, and $m_{0}(\mathcal{F}_{m})=m+2$. The reduction of singularities of the pair $(\mathcal{F}_{m},B)$ is achieved in $2m$ steps, and $B\in\text{Dic}_{0}(\mathcal{F}_{m})$. Thus
$\text{DG}_{0}(\mathcal{F}_{1},B)$ $=$ $\left[\nu_{0}(B)(2\nu_{0}(B)-m_{0}(\mathcal{F})-1)-4\alpha_{0}\right]$
$+\left[\nu_{q^{(1)}}(B^{(1)})(2\nu_{q^{(1)}}(B^{(1)})-m_{q^{(1)}}(\mathcal{F}^{(1)})-1)-4\alpha_{1}\right]$
$=$ $\left[2(4-3-1)-4(1)\right]+\left[1(2-2-1)-4(0)\right]$
$=$ $-5$
and for $m>1$, we have
$\text{DG}_{0}(\mathcal{F}_{m},B)$ $=$ $\sum_{j=0}^{2m-1}\left[\nu_{q^{(j)}}(B^{(j)})(2\nu_{q^{(j)}}(B^{(j)})-m_{q^{(j)}}(\mathcal{F}^{(j)}_{m})-1)-4\alpha_{j}\right]$
$=$ $\left[\nu_{0}(B)(2\nu_{0}(B)-m_{0}(\mathcal{F}_{2m})-1)-4\alpha_{0}\right]$
$+\left[\nu_{q^{(2m-1)}}(B^{(2m-1)})(2\nu_{q^{(2m-1)}}(B^{(2m-1)})-m_{q^{(2m-1)}}(\mathcal{F}^{(2m-1)}_{2m})-1)-4\alpha_{2m-1}\right]$
$=$ $\left[2m\left(2(2m)-(m+2)-1\right)-4(m^{2}+m)\right]+\left[1(2-2-1)-4(0)\right]$
$=$ $1\displaystyle 2m^{2}-10m-$ 1 .
Therefore, $\text{DG}_{0}(\mathcal{F}_{m},B)\geq 0$ if and only if $m\geq 6$. This shows that the hypothesis of Theorem A is optimal.
5. Applications
In this section, from the proof of Theorem A we deduce the Dimca-Greuel bound for the case of irreducible plane curve germs, proved simultaneously in [1], [11] and [19].
@@ 5. Applications
Let $f(x,y)\in\mathbb{C}\{x,y\}$ an irreducible power series with an isolated singularity at $0\in\mathbb{C}^{2}$. Put $C:f(x,y)=0$ the branch defined by $f$. Then
@@
Consider the singular holomorphic foliation $\mathcal{F}:\omega=df$. In this situation, $C\in\text{ISO}_{0}(\mathcal{F})$. To continue, we compute
$\text{DG}_{0}(\mathcal{F},C)=R_{0}(\mathcal{F},C)-\tau_{q^{(N)}}(\mathcal{F}^{(N)},C^{(N)}),$
where
$R_{0}(\mathcal{F},C):=\sum_{j=0}^{N-1}\left[\nu_{q^{(j)}}(C^{(j)})(2\nu_{q^{(j)}}(C^{(j)})-m_{q^{(j)}}(\mathcal{F}^{(j)})-1)-4\alpha_{j}\right],$
and $C^{(j+1)}$ denotes the strict transform of $C^{(j)}$ thought $q^{(j)}$ for each $j=0,\ldots,N-1$, $C^{(N)}$ is the normalization of $C$ (or a strict transform of it), $\mathcal{F}^{(i+1)}$ denotes the strict transform of $\mathcal{F}^{(j)}$, and $\mathcal{F}^{(N)}$ is a foliation with a reduced singularity at $q^{(N)}\in C^{(N)}$ as in the proof of Theorem A.
Since $C^{(N)}$ is smooth, we have $\tau_{q^{(N)}}(\mathcal{F}^{(N)},C^{(N)})=GSV_{q^{(N)}}(\mathcal{F}^{(N)},C^{(N)})$ by Remark 4.2, and from Lemma 3.1, we deduce
$\text{DG}_{0}(\mathcal{F},C)=-GSV_{0}(\mathcal{F},C)+\sum_{j=0}^{N-1}\left[\nu_{q^{(j)}}(C^{(j)})(\nu_{q^{(j)}}(C^{(j)})-1)-4\alpha_{j}\right]$
Now, since $\mathcal{F}$ is a curve generalized non-dicritical foliation at $0\in\mathbb{C}^{2}$, and $\text{Sep}_{0}(\mathcal{F})=C$, we have $GSV_{0}(\mathcal{F},C)=0$ by [3, p. 528]. Moreover $\mu_{0}(\mathcal{F},C)=\mu(C)$ and $\tau_{0}(\mathcal{F},C)=\tau(C)$ by Remark 2.1. Hence
$\text{DG}_{0}(\mathcal{F},C)=\sum_{j=0}^{N-1}\left[\nu_{q^{(j)}}(C^{(j)})(\nu_{q^{(j)}}(C^{(j)})-1)-4\alpha_{j}\right].$
Finally, applying successively [19, Claim 4.4] in each step of the resolution of singularities of $C$, and given that the corresponding value $\mathcal{D}_{min}$ of [19, Claim 4.4] in the $j$-th step of the resolution of $C$ is less or equal to $\alpha_{j}$, we get $\text{DG}_{0}(\mathcal{F},C)<0$ implying
$\frac{\mu(C)}{\tau(C)}<\frac{4}{3}$
by Theorem A. ∎
Now, if in (1), we change $\mu_{p}(\mathcal{F})$ to the multiplicity of the singular foliation $\mathcal{F}$ along a reduced curve $C$, $\mu_{p}(\mathcal{F},C)$ (see Section 2), and we put in the denominator of (1) the sum $T_{p}(\mathcal{F},C):=\sum_{B}\tau_{p}(\mathcal{F},B)$ of Tjurina numbers of $\mathcal{F}$ along the irreducible components of $C$, then we get the following inequality of the Dimca-Greuel type for foliations:
@@
Let $\mathcal{F}$ be a germ of a singular holomorphic foliation at $(\mathbb{C}^{2},p)$. Let $\mathcal{B}$ be an effective primitive balanced divisor of separatrices for $\mathcal{F}$ at $p$. If $\mathcal{B}=\sum_{j=1}^{\ell}B_{j}$ with $\text{DG}_{p}(\mathcal{F},B_{j})<0$ for all $1\leq j\leq\ell$, then
(28) $\frac{\mu_{p}(\mathcal{F},\mathcal{B})}{T_{p}(\mathcal{F},\mathcal{B})}<\frac{4}{3}.$
@@
Suppose that $\mathcal{B}=\sum_{j=1}^{\ell}B_{j}$, where each $B_{j}$ is a separatrix for $\mathcal{F}$. It follows from Theorem A that
@@
$\frac{\mu_{p}(\mathcal{F},\mathcal{B})}{T_{p}(\mathcal{F},\mathcal{B})-\ell+1}<\frac{4}{3}.$
Since $\ell\geq 1$, we obtain $\frac{\mu_{p}(\mathcal{F},\mathcal{B})}{T_{p}(\mathcal{F},\mathcal{B})}\leq\frac{\mu_{p}(\mathcal{F},\mathcal{B})}{T_{p}(\mathcal{F},\mathcal{B})-\ell+1}<\frac{4}{3}.$